# Deep Support Vectors

### v1

@@ 5.2 DSVs: Decoding Black-Box in Deep Learning
In Fig. 5, we conducted an experiment by mixing a randomly sampled image from the real dataset with an image from the DSVs. Upon observation, the mixed image is virtually indistinguishable from an image obtained solely from the real dataset. It is noteworthy to highlight this situation resembles that of an adversarial attack (Xu et al., 2019; Moosavi-Dezfooli et al., 2016), yet we did not apply gradient descent to the image; we simply mixed two images. This suggests that the discriminative informational density in a single DSV image is substantially greater than that in a randomly sampled image. The fact that the DSV’s characteristics remained dominant in the classification, underscores the significant role of DSVs in explaining the model’s classification ability.
Figure 6: Model predictions for original versus DSV-informed edited images. Images were altered based on decision criteria derived from DSVs, influencing the model’s prediction.
DSVs Encode the Decision Criteria Visually
Our findings suggest that DSVs for image classification problems not only satisfy the conditions of classical support vectors but also encapsulate a greater volume of visual information. Figure 6 experimentally substantiates our claims and illustrates the practical use of DSVs. Analysis of Figure 1 reveals the decision criteria used for classifying deer, cats, and dogs: 1) DSVs highlight antlers in deer, signifying them as a distinctive characteristic. 2) In cats, pointed triangular ears are a recurring feature in DSVs. 3) For dogs, a trio of facial dots holds significant importance. Using these observations, we altered a deer’s image by erasing its antlers, thereby reducing the model’s confidence in classifying it as a deer, and by reshaping its ears to a pointed contour, which caused the model to misclassify it as a cat. Similarly, by smoothing the ears of a cat image to diminish its classification confidence and then adding antlers or three facial dots, we influenced the model to reclassify the image as a deer or a dog, respectively, according to the introduced features.
DSVs Enable Qualitative Assessment of a Model
This discovery holds significant implications about making models responsible; it introduces a qualitative aspect to assessing model performance. Consider a scenario where we need to differentiate between deer and horses, and the data is collected in summer. The model in our study would be less suitable, as evidenced by Fig. 1, which indicates the model’s reliance on antlers for identifying deer – a feature not present in summer. This shows that DSVs enable us to conduct causal predictions by qualitatively analyzing models, an aspect integral to RAI (Ghallab, 2019). Without DSVs, reliance would be solely on scoring criteria, a potentially hazardous dependence given that model performance is heavily reliant on the quality of data. If the data is biased or contaminated, the model’s ability to make accurate judgments could be compromised. DSVs help mitigate such risks, providing a clearer and more responsible understanding of the model’s decision-making processes.
DSVs: Reconstructing Models With a Minimalist Touch
SVMs utilize support vectors to construct classification models, and this research focuses on DSVs, which serve a similar role to SVM’s support vectors. We have established that DSVs meet the theoretical conditions required by SVM’s support vectors. Based on this, we explored the possibility of reconstructing models using DSVs.
In Table 1, we demonstrate that DSVs can be effectively used to retrain models with a limited amount of training data, aligning with the principles of dataset distillation. Specifically, our experiments were conducted in two approaches, as shown in Fig. 1: coreset selection and image synthesis. For each class, we either generated or selected a single image and these images were used to retrain models from scratch. In terms of image selection, our approach outperformed naive random selection and the forgetting algorithm. This implies that our selection process itself is effective. In the realm of image synthesis, our technique achieved results on par with established dataset distillation methods, surpassing the coreset selection approaches. These findings are noteworthy, especially considering that traditional distillation algorithms depend on strong conditions that they can access to the entire dataset, and have multiple trained models with different initializations. In contrast, our algorithm operates relying solely on a single trained model and does not require any data during the generation process, not even any knowledge about prior dataset as referenced in (Haim et al., 2022). This suggests that the ability of our approach to match the performance of traditional methods, despite not directly utilizing a loss function to enhance the accuracy of retrained models, is a strong indication of the validity of our assertion. It implies that DSVs inherently possess characteristics similar to support vectors in SVMs, as only this could explain their effectiveness in the absence of the direct optimization strategies employed by conventional algorithms.
Table 1: An experimental result of datastet distillation. DD refers to the dataset distillation algorithm in (Wang et al., 2018) and DC refers to the dataset condensation algorithm in (Zhao et al., 2020). In coreset selection, herding refers to  (Chen et al., 2012) and forgetting refers to  (Toneva et al., 2018). DSVs and DSVg represent our method of DSV selection and DSV generation respectively.
coreset selection synthesis
random herding forgetting DSVs DD DC DSVg
14.4 $\pm$ 2.0 21.2 $\pm$ 1.2 13.3 $\pm$ 1.2 17.45 $\pm$ 1.2 27.3 $\pm$ 0.8 28.3 $\pm$ 0.5 25.3 $\pm$ 0.5
DSVs, the First Practical Dataset Distillation
In the context of dataset distillation, DeepKKT can be used as a pioneering algorithm tailored for practical scenarios. Dataset distillation primarily addresses two key issues: 1) Protecting private information through data synthesis, and 2) Reducing the communication load by minimizing the data size. Traditional algorithms, however, face a paradox, particularly when data originates predominantly from edge devices like smartphones (McMahan et al., 2016; Konečný et al., 2016). These algorithms typically require access to the entire dataset, presumably centralized, leading to significant communication overhead and privacy concerns. Contrasting with these approaches, DeepKKT operates independently of any dataset, relying solely on a pre-trained model. This approach eliminates the need for edge devices to store extensive private data. Additionally, due to its lightweight nature, DSVs can be effectively extracted directly on edge devices.
Figure 7: Images of DSVs generated by a model that underwent transfer learning.
DSVs in Practical Settings
[…]
6 Ablation study
Effect of Each Loss
In the synthesis of DSVs, the process starts from noise, which is sampled from a Gaussian distribution. While this initial step resembles the diffusion process in diffusion models (Ho et al., 2020; Song & Ermon, 2019), there is a fundamental difference in the target of the transition. The only difference is its desired distribution. In diffusion models, the transition aims to capture the entire data manifold. However, for DSV synthesis, the focus is specifically on $\partial\mathcal{M}$, the boundary of the data manifold.
A discretized diffusion model with classifier guidance can be described by the following equation:
$x_{t+1}=x_{t}+\epsilon_{t}\cdot\left(\nabla_{x}\log p(x_{t})+\lambda\nabla_{x}\log p(y|x_{t})\right)$ (9)
Our Algorithm 1, which can be formulated as:
$x_{t+1}=x_{t}+\alpha\cdot\left(\nabla_{x}L_{\text{stat}}(A(x))+\beta\nabla_{x}L_{\text{primal}}(A(x)\right)$ (10)
Where $A$ is random augmentation function. In our algorithm, the stationarity condition mirrors the score in diffusion models, while the primal condition aligns with the concept of guidance. However, unlike traditional diffusion models that target the whole data manifold, our process is directed towards navigating from the Gaussian distribution to $\partial\mathcal{M}$, the boundary of the data manifold. This focus on the manifold boundary is a distinctive aspect of our approach, emphasizing the generation of DSVs within the crucial decision boundaries. Fig. 8 also supports our hypothesis that the primal condition alone does not generate meaningful images.. Moreover, while stationarity loss alone can create images with well-structured shapes, it fails to generate accurate samples if the sample labels do not match the targeted label, especially in the absence of primal guidance (classifier guidance). Exploring the intricate relationship between stationarity and diffusion in the context of data manifold navigation presents an exciting avenue for future research.
Figure 8: Results showing DeepKKT images created solely by the primal condition or by the stationary condition.
7 Conclusion
This paper pioneers the definition of support vectors in nonlinear deep learning models, specifically through the introduction of DSVs (Deep Support Vectors). We demonstrated the feasibility of generating DSVs using only pretrained models, without requiring access to the training dataset. To achieve this, we extended the KKT (Karush-Kuhn-Tucker) conditions, to DeepKKT conditions. It can be applied to any deep learning models. The synthesized DSVs effectively encode the decision boundary, offering a visual and intuitive explanation of a model’s decision-making process. Furthermore, we have shown that models can be retrained using DSVs in a manner akin to SVMs.
Societal Impacts
Our paper is closely related to Responsible AI (RAI), particularly in terms of enabling qualitative assessment of models. They provide a visual and intuitive explanation of a model’s decision-making criteria, offering insights into how the model operates that are not only explanatory but also responsible. This approach goes beyond traditional Explainable AI (XAI) methodologies, which typically focus on local explanations for individual inputs and lack the ability to offer a global decision criterion like DSVs. Additionally, as this method is based on model inversion, it ensures safety and privacy. Although the synthesized sets in Fig. 1 might look similar to the selected sets, they do not replicate specific sample features. This is because DSVs represent a more generalized decision boundary, thus avoiding the inclusion of image-specific features. Therefore, DSVs can make all models using logistic loss responsible.
Appendix A intutive explanation of DeepKKT condition
[…]
In DeepKKT, many conditions, except one condition makes sense. For example, primal condition and manifold condition makes sense and dual condition can be regraded as importance sampling. The most counterinutitive part is stationarity section.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}$ (11)
In this section, we’re to explain the dynamics of overparameterized deep learning, and explain how it is connected to deep learning. Below is an analogy of (Soudry et al., 2018).
A deep learning models follows following ODE:
$w_{t+1}=w_{t}-\alpha\nabla L(x,y;w_{t})$ (12)
Where $\alpha$ is the learning rate and $t$ is the optimization step. $L$ does not go to zero since deep learning models usually exploit a function with a logistic tail, such as the cross-entropy function. Thus, there exists a convergence of the gradient direction $g_{\infty}\coloneqq\hat{\nabla}L$. There also exists a time $T$ where the gradient direction converges to $g_{\infty}-\varepsilon$ for sufficiently small $\varepsilon$. As illustrated in Fig. 9, $w$ moves toward the direction of $-g^{\infty}$. Therefore, $\hat{w}_{\infty}\approx-g_{\infty}$.
This is what stationarity condition wants to seek for. The direction of $g_{\infty}$, by using only few support vectors.
Appendix B Full Implementation Details
For $L_{\text{stat}}$, we set $\alpha_{x}=10^{-3}$, $\alpha_{\lambda}=0.1$, and $\gamma=5$, respectively. This was due to the differing scales. The actual computed loss for stationarity was ten times higher than for primal loss. To satisfy the manifold condition in reconstructing DSVs, we implemented techniques from the diffusion model (Rombach et al., 2021). This involved synthesizing data at a low resolution, then upscaling it to a higher resolution. We initially downscaled the image twice, then upscaled it, continuing with optimization. To reduce the impact of outliers, we choose the $L_{1}$ metric and apply the square root to the gradient term when measuring distance.
To obtain the pretrained weight $\theta^{*}$, we chose the ConvNet architecture (Goodfellow et al., 2016), a common choice in deep learning. This architecture includes sequential convolutional layers followed by max pooling, and a single fully-connected layer for classification. The learning rate was set to $10^{-3}$ with a weight decay of $0.005$ using the Adam optimizer. Additionally, we employed flipping and cropping techniques, with settings differing from those used for DSVs reconstruction to ensure fair comparison. For pretraining $\Phi$ on the Street View House Numbers (SVHN) dataset (Netzer et al., 2011), a digit dataset with dimensions similar to CIFAR-10 (Krizhevsky, 2009), we exclusively trained the fully-connected layer of the CIFAR-10 pre-trained ConvNet. This approach resulted in a training accuracy of 80%. When retraining the model with DSVs, we used large learning rate of 0.1 and a $\rho$ value of 0.0001 were utilized for the SAM optimizer to adhere to the principle of margin maximization (Wang et al., 2023; Foret et al., 2020). For DSVs synthesis, we used a single 1080TI GPU and it took 15 minutes.
Figure 10: Images of DSV candidates (Selection).
Figure 11: Images of various DSVs (Synthesized) – Cat, Ship, Dog, Car, Truck, Bird, Horse, Deer, Plane, Frog.
Figure 12: Synthesized images on other architecture.

### v2

@@ Appendix F More Characteristics of DSVs
In Fig. 9, we conducted an experiment by mixing a randomly sampled image from the real dataset with an image from the DSVs. Upon observation, the mixed image is virtually indistinguishable from an image obtained solely from the real dataset. It is noteworthy to highlight this situation resembles that of an adversarial attack [32, 19], yet we did not apply gradient descent to the image; we simply mixed two images. This suggests that the discriminative informational density in a single DSV image is substantially greater than that in a randomly sampled image. The fact that the DSV’s characteristics remained dominant in the classification, underscores the significant role of DSVs in explaining the model’s classification ability.
Appendix G More Examples
In Fig. 10, examples of latent interpolations between target labels are presented. The smoothness of these interpolations within the latent space indicates that the semantic information learned from the training data has been effectively applied during the DSV generation process. This observation provides evidence that the DeepKKT optimization successfully conducts the generative process.
Fig. 11 and  12 provide examples of deep support vectors generated using the CIFAR-100 and ImageNet datasets, respectively. Fig. 13 presents additional examples related to image editing.
Fig. 14 empirically supports on our assertion on decision criterion. Starting from CIFAR100 random images and CIFAR10-pretrained models, we edited the image CIFAR10 labels as latents. The edited images changes the image following decision criterions in generated DSVs. 1) For editing images to deer, antler grows. 2) For dog editing, facial dots are generated. 3) For cat editing, pointed traiangler features are generated.
Appendix H Algorithm
Alg. 1 presents our algorithm of generating Deep Support Vectors (DSVs). Initialized either from a noise $x_{i}^{s}\sim\mathcal{N}(0,I)$ or a real sample, it iterates to obtain the primal $X^{S}$ and dual $\Lambda^{S}$ variables.
Algorithm 1 Support Vector Refinement for Deep Learning Model
1:Pretrained classifier $\Phi(\cdot;\theta)$, loss function $L$, augmentation function set $A$, number of DSV candidate $N$, number of class $C$, hyperparameters $\alpha$, $\beta$
2:Freeze classifier $\Phi(\cdot;\theta)$
3:Initialize $N\times C$ number of support vector candidates
4:for $i=1$ to $C$ do
5:     sample $N$ number of $(x^{s}_{i},\lambda^{s}_{i})$ for label $y^{s}_{i}=i$
6:end for
7:Define $X^{S}=\{x^{s}_{i}\mid i\in[C],s\in[N]\}$
8:Define $\Lambda^{S}=\{\lambda^{s}_{i}\mid i\in[C],s\in[N]\}$
9:repeat
10:     $L_{\text{primal}}(X^{S})=\sum_{s=1}^{N}\sum_{i=1}^{C}L(\Phi(x^{s}_{i};\theta),y^{s}_{i})$
11:     $L_{\text{stationary}}(X^{S})=\|\theta+\sum_{s=1}^{N}\sum_{i=1}^{C}\lambda^{s}_{i}y^{s}_{i}\nabla_{\theta}\Phi(x^{s}_{i};\theta)\|^{2}_{2}$
12:     $L_{\text{kkt}}(X^{S})=\beta_{1}\cdot L_{\text{primal}}(X^{S})+L_{\text{stationary}}(X^{S})$
13:     $L_{\text{prior}}$ = $\beta_{2}\cdot L_{\text{tot}}(X)+\beta_{3}L_{\text{norm}}(X)$
14:     Sample $f_{A}\in A$
15:     Define $AX^{S}=\{f_{A}(x^{s}_{i})\mid x^{s}_{i}\in X^{S}\}$
16:     $L_{\text{akkt}}(X^{S})=L_{\text{kkt}}(AX^{S})$
17:     $L_{\text{total}}(X^{S})=L_{\text{kkt}}(X^{S})+\eta\cdot L_{\text{akkt}}(X^{S})$ + $L_{\text{prior}}$
18:     Update $X^{S}\leftarrow X^{S}+\nabla_{X^{S}}L_{\text{total}}(X^{S})$
19:     Update $\Lambda^{S}\leftarrow\Lambda^{S}+\nabla_{\Lambda^{S}}L_{\text{total}}(X^{S})$
20:     Remove $x^{s}_{i}$s for corresponding $\lambda^{s}_{i}<0$
21:until $X^{S}$ converges
22:return Set of DSV : $X^{S}$